# Sparse Autoencoders for Scientifically Rigorous Interpretation of Vision Models

### v1

@@ Appendix A SAE Training Details
We then train a newly initialized SAE on $100$M activations. $W_{\text{enc}}$ and $W_{\text{dec}}$ are initialized using PyTorch’s kaiming_uniform_ initialization and $b_{\text{enc}}$ is zero-initialized. $b_{\text{dec}}$ is initialized as the mean of $524,288$ random samples from the dataset, as recommended by prior work (Templeton et al., 2024b; a).
We use mean squared error $||\mathbf{\hat{x}}-\mathbf{x}||_{2}^{2}$ as our reconstruction loss, L1 length of $f(\mathbf{x})$ scaled by the current sparsity coefficient $\lambda$ as our sparsity loss, and track L0 throughout training.
The learning rate $\eta$ and sparsity coefficient $\lambda$ are linearly scaled from 00 to their maximum over 500 steps each and remain at their maximum for the duration of training.
@@
Hidden Width $24,576$ (32×32$\times$32 × expansion)
@@
Table A1: Hyperparameters for training SAEs. Sparsity coefficient $\lambda$ and learning rate $\eta$ are chosen qualitatively by inspecting discovered features.
Appendix B Additional Examples
Additional examples of CLIP learning cultural features are in Figure B1.
@@ Appendix B Additional Examples
Figure B1: Additional examples of cultural features learned by CLIP. Top: CLIP-24K/7622 responds to symbolism from the United States of America, including a portrait of George Washington, but not to a portrait of King Louis XIV of France. Bottom: CLIP-24K/13871 activates strongly on the Brandenburg Gate and other German symbols, but not on visually similar flags like the Belgian flag.
Appendix C Classification Details
We use an SAE trained on $100$M CLIP ViT-B/$16$ activations from layer $11$ on ImageNet-1K’s training split of $1.2$M images following the procedure in Appendix A.
C.1 Task-Specific Decoder Training (Image Classification)
For image classification, we train a linear probe on the [CLS] token representations extracted from a CLIP-pretrained ViT-B/16 model (Radford et al., 2021) using the CUB-2011 dataset (Wah et al., 2011). The following details describe our experimental setup to ensure exact reproducibility.
@@ C.1 Task-Specific Decoder Training (Image Classification)
Each input image is first resized so that its shorter side is 256 pixels, followed by a center crop to obtain a 224$\times$224 image. The resulting image is then normalized using the standard ImageNet mean and standard deviation for RGB channels. No additional data augmentation is applied.
@@
We train the linear probe on the CUB-2011 training set using the AdamW (Loshchilov, 2017) optimizer for 20 epochs with a batch size of 512. We performed hyperparameter sweeps over the learning rate with values in {$10^{-5}$, $10^{-4}$, $10^{-3}$, $10^{-2}$, $10^{-1}$, $1.0$} and over weight decay with values in {$0.1$, $0.3$, $1.0$, $3.0$, $10.0$}. Based on validation accuracy, we selected a learning rate of $10^{-3}$ and a weight decay of 0.1. The CLIP backbone remains frozen during this process. The trained linear probe achieves a final accuracy of $79.9$% on the CUB-2011 validation set.
C.2 Additional Examples
We provide additional examples of observing classification predictions, using SAEs to form a hypothesis explaining model behavior, and experimentally validating said hypothesis through feature suppression in Figures C2, C3 and C4. Each figure links to a live, web-based demo where readers can complete the “observe, hypothesize, experiment” sequence themselves.
Figure C2: Tropical Kingbirds have a distinctive yellow chest. When we suppress a “yellow feathers” feature, our linear classifier predicts Gray Kingbird, a similar species but with a gray chest. This example is available at https://osu-nlp-group.github.io/SAE-V/demos/classification?example=5099
Figure C3: Canada Warblers have a distinctive black necklace on the chest. CLIP-24K/20376 fires on similar patterns; when we suppress this feature, the linear classifier predicts Wilson Warbler, a similar species without the distinctive black necklace. This example is available at https://osu-nlp-group.github.io/SAE-V/demos/classification?example=1129
Figure C4: Purple finches have bright red coloration on the head and neck area; when we suppress CLIP-24K/10273, which appears to be a “red feathers” feature, our classifier predicts Field Sparrow, which has similar wing banding but no red coloration. This example is available at https://osu-nlp-group.github.io/SAE-V/demos/classification?example=4139
Appendix D Semantic Segmentation Details
@@
D.1 Task-Specific Decoder Training (Semantic Segmentation)
For semantic segmentation, we train a single linear segmentation head on features extracted from a frozen DINOv2-pretrained ViT-B/14 model (Oquab et al., 2023) using the ADE20K dataset (Zhou et al., 2017). We aim to map patch-level features to 151 semantic class logits. The following details describe our experimental setup to ensure exact reproducibility.
@@ D.1 Task-Specific Decoder Training (Semantic Segmentation)
Each image is first resized so that its shorter side is 256 pixels, followed by a center crop to obtain a 224$\times$224 image. The cropped images are normalized using the standard ImageNet mean and standard deviation for RGB channels. No additional data augmentation is applied.
@@
The segmentation head is a linear layer applied independently to each patch token. This layer maps the $d$-dimensional feature vector to a $151$-dimensional logit vector, corresponding to the $151$ semantic classes. For visualization purposes, we perform a simple upsampling by replicating each patch prediction to cover the corresponding $14\times 14$ pixel block. For quantitative evaluation (computing mIoU), the $16\times 16$ logit map is bilinearly interpolated to the full image resolution.
@@
We train the segmentation head using the standard cross entropy loss. The DINOv2 ViT-B/14 backbone remains frozen during training, so that only the segmentation head is updated. We use AdamW (Loshchilov, 2017) with a batch size of $1,024$ over $400$ epochs. We performed hyperparameter sweeps over the learning rate with values in $3\times 10^{-5}$, $1\times 10^{-4}$, $3\times 10^{-4}$, $1\times 10^{-3}$, $3\times 10^{-3}$, $1\times 10^{-2}$} and over weight decay values in $1\times 10^{-1}$, $1\times 10^{-3}$, $1\times 10^{-5}$}. No learning rate schedule is applied; the chosen learning rate is kept constant throughout training. The segmentation head is initialized with PyTorch’s default initialization.
@@
D.2 Additional Examples
We provide additional examples of observing segmentation predictions, using SAEs to form a hypothesis explaining model behavior, and experimentally validating said hypothesis through feature suppression in Figures D5, D6, D7 and D8. These additional examples further support the idea of SAEs discovering features that are pseduo-orthogonal. Each figure links to a live, web-based demo where readers can complete the “observe, hypothesize, experiment” sequence themselves.
Figure D5: DINOv2 correctly identifies the framed painting. We find that DINOv2-24K/16446 fires for paintings, and that suppressing this feature removes the painting without meaningfully affecting other parts of the image, despite modifying all patches. This example is available at https://osu-nlp-group.github.io/SAE-V/demos/semseg?example=1633.
Figure D6: We find that feature DINOv2-24K/5876 fires for toilets, and that suppressing this feature removes the toilet without meaningfully affecting other parts of the image, despite modifying all patches. This example is available at https://osu-nlp-group.github.io/SAE-V/demos/semseg?example=1099.
Figure D7: After suppressing a “bed” feature (DINOv2-24K/18834), the segmentation head predicts “pillow” for the pillows and “table” for the bed spread. This examples is available at https://osu-nlp-group.github.io/SAE-V/demos/semseg?example=1117.
Figure D8: We remove all cars from the scene by suppressing a “car”-like feature (DINOv2-24K/7235). Notably, the van on the right of the image is also removed. This examples is available at https://osu-nlp-group.github.io/SAE-V/demos/semseg?example=1852.

### v2

@@ 8 SAE Training Details
We then train a newly initialized SAE on 100100M activations. $W_{\text{enc}}$ and $W_{\text{dec}}$ are initialized using PyTorch’s kaiming_uniform_ initialization and $b_{\text{enc}}$ is zero-initialized. $b_{\text{dec}}$ is initialized as the mean of $524,288$ random samples from the dataset, as recommended by prior work [templeton2024scaling, templeton2024update].
We use mean squared error $\mathbf{\hat{x}}-\mathbf{x}||_{2}^{2}$ as our reconstruction loss, L1 length of $f(\mathbf{x})$ scaled by the current sparsity coefficient $\lambda$ as our sparsity loss, and track L0 throughout training.
The learning rate $\eta$ and sparsity coefficient $\lambda$ are linearly scaled from 0 to their maximum over 500 steps each and remain at their maximum for the duration of training.
@@
Table 2: Hyperparameters for training SAEs. Sparsity coefficient $\lambda$ and learning rate $\eta$ are chosen qualitatively by inspecting discovered features.
@@
Hidden Width $24,576$ ($32\times$ expansion)
@@
9 SAE-Enabled Analysis Details
For our comparative analysis of CLIP and DINOv2 using sparse autoencoders (SAEs), we employ a qualitative discovery approach focused on identifying meaningful semantic features learned by each model. This section details our experimental methodology.
Model and Layer Selection:
We use pretrained CLIP ViT-B/16 [radford2021learning] and DINOv2 ViT-B/14 [oquab2023dinov2] models as our base vision transformers. For both models, we extract activations from the $11$th transformer layer (second-to-last layer), as features at this depth capture high-level semantic information while not being overly specialized to the model’s specific pretraining objective.
SAE Architecture and Training:
Following the procedure detailed in Section˜3, we train identical SAE architectures for both models. We use the same preprocessing and normalization steps as described in the main methodology section, ensuring consistency across experiments.
Feature Discovery Process:
Our feature discovery process uses a manual exploration approach:
  1. 1.
    We build a custom UI for feature visualization that displays exemplar images that maximally activate SAE features.
  2. 2.
    Features are sorted by two criteria: sparsity (lower is better) and maximum activation value (higher is better).
  3. 3.
    We systematically examine hundreds of features, recording observations about semantic patterns.
  4. 4.
    When we find potentially interesting features (e.g., country-specific representations or abstract concepts), we conduct further targeted investigations.
This iterative process involves multiple hours of exploration per model to identify prominent patterns in feature representations between CLIP and DINOv2.
Feature Verification:
To verify the semantic consistency of discovered features, we employ a deliberate testing approach:
  1. 1.
    For each potentially interesting feature, we examine both positive examples (images expected to contain the concept) and negative examples (images without the concept).
  2. 2.
    We deliberately select challenging negative examples that shared visual similarities but differed in the target semantic concept (e.g., national symbols from different countries, visually similar but semantically different accident scenes).
  3. 3.
    For cross-model comparison, we attempt to find corresponding features in both models by:
    • Starting with positive examples known to activate a specific feature in one model.
    • Identifying features in the second model with high activation on the same examples.
    • Examining maximally activating examples for those candidate features to verify semantic alignment.
Exemplar Selection for Figures:
We manually select exemplar images in figures to effectively demonstrate the semantic properties of each feature. We choose positive examples to illustrate the breadth of a concept (e.g., different manifestations of “Brazil” or “accidents”), and negative examples to demonstrate feature specificity (e.g., other South American imagery that does not activate the “Brazil” feature). Additional examples of CLIP learning cultural features are in Fig.˜8.
@@ 9 SAE-Enabled Analysis Details
Figure 8: Additional examples of cultural features learned by CLIP. Top: CLIP-24K/7622 responds to symbolism from the United States of America, including a portrait of George Washington, but not to a portrait of King Louis XIV of France. Bottom: CLIP-24K/13871 activates strongly on the Brandenburg Gate and other German symbols, but not on visually similar flags like the Belgian flag.
10 Classification Details
We use an SAE trained on 100100M CLIP ViT-B/$16$ activations from layer $11$ on ImageNet-1K’s training split of 1.21.2M images following the procedure in Section˜8.
10.1 Task-Specific Decoder Training (Image Classification)
For image classification, we train a linear probe on the [CLS] token representations extracted from a CLIP-pretrained ViT-B/16 model [radford2021learning] using the CUB-2011 dataset [wah2011cub]. The following details describe our experimental setup to ensure exact reproducibility.
@@ 10.1 Task-Specific Decoder Training (Image Classification)
Each input image is first resized so that its shorter side is 256 pixels, followed by a center crop to obtain a $\times224$ image. The resulting image is then normalized using the standard ImageNet mean and standard deviation for RGB channels. No additional data augmentation is applied.
@@
We train the linear probe on the CUB-2011 training set using the AdamW [loshchilov2019adamw] optimizer for 20 epochs with a batch size of 512. We performed hyperparameter sweeps over the learning rate with values in $10^{-5}$, $10^{-4}$, $10^{-3}$, $10^{-2}$, $^{-1}, 1.01.0$} and over weight decay with values in {$0.1$, $0.3$, $1.0$, $3.0$, 10.010.0}. Based on validation accuracy, we selected a learning rate of $10^{-3}$ and a weight decay of 0.1. The CLIP backbone remains frozen during this process. The trained linear probe achieves a final accuracy of 79.979.9% on the CUB-2011 validation set.
10.2 Additional Examples
We provide additional examples of observing classification predictions, using SAEs to form a hypothesis explaining model behavior, and experimentally validating said hypothesis through feature suppression in Figs.˜9, 10 and 11. Each figure links to a live, web-based demo where readers can complete the “observe, hypothesize, experiment” sequence themselves.
Figure 9: Tropical Kingbirds have a distinctive yellow chest. When we suppress a “yellow feathers” feature, our linear classifier predicts Gray Kingbird, a similar species but with a gray chest. This example is available at https://anonymous-saev.github.io/saev/demos/classification?example=5099
Figure 10: Canada Warblers have a distinctive black necklace on the chest. CLIP-24K/20376 fires on similar patterns; when we suppress this feature, the linear classifier predicts Wilson Warbler, a similar species without the distinctive black necklace. This example is available at https://anonymous-saev.github.io/saev/demos/classification?example=1129
Figure 11: Purple finches have bright red coloration on the head and neck area; when we suppress CLIP-24K/10273, which appears to be a “red feathers” feature, our classifier predicts Field Sparrow, which has similar wing banding but no red coloration. This example is available at https://anonymous-saev.github.io/saev/demos/classification?example=4139
11 Semantic Segmentation Details
@@
11.1 Task-Specific Decoder Training (Semantic Segmentation)
For semantic segmentation, we train a single linear segmentation head on features extracted from a frozen DINOv2-pretrained ViT-B/14 model [oquab2023dinov2] using the ADE20K dataset [zhou2017ade20k]. We aim to map patch-level features to 150 semantic class logits. The following details describe our experimental setup to ensure exact reproducibility.
@@ 11.1 Task-Specific Decoder Training (Semantic Segmentation)
Each image is first resized so that its shorter side is 256 pixels, followed by a center crop to obtain a $\times224$ image. The cropped images are normalized using the standard ImageNet mean and standard deviation for RGB channels. No additional data augmentation is applied.
@@
The segmentation head is a linear layer applied independently to each patch token. This layer maps the $d$-dimensional feature vector to a $150$-dimensional logit vector, corresponding to the $150$ semantic classes. For visualization purposes, we perform a simple upsampling by replicating each patch prediction to cover the corresponding $14\times 14$ pixel block. For quantitative evaluation (computing mIoU), the $16\times 16$ logit map is bilinearly interpolated to the full image resolution.
@@
We train the segmentation head using the standard cross entropy loss. The DINOv2 ViT-B/14 backbone remains frozen during training, so that only the segmentation head is updated. We use AdamW [loshchilov2019adamw] with a batch size of $1,024$ over $400$ epochs. We performed hyperparameter sweeps over the learning rate with values in $3\times 10^{-5}$, $1\times 10^{-4}$, $3\times 10^{-4}$, $1\times 10^{-3}$, $3\times 10^{-3}$, $1\times 10^{-2}$} and over weight decay values in $\times 10^{-1}, 1×1$$\times 10^{-3}$, $1\times 10^{-5}$}. No learning rate schedule is applied; the chosen learning rate is kept constant throughout training. The segmentation head is initialized with PyTorch’s default initialization.
@@
11.2 Additional Examples
We provide additional examples of observing segmentation predictions, using SAEs to form a hypothesis explaining model behavior, and experimentally validating said hypothesis through feature suppression in Figs.˜12, 13, 14 and 15. These additional examples further support the idea of SAEs discovering features that are pseduo-orthogonal. Each figure links to a live, web-based demo where readers can complete the “observe, hypothesize, experiment” sequence themselves.
Figure 12: DINOv2 correctly identifies the framed painting. We find that DINOv2-24K/16446 fires for paintings, and that suppressing this feature removes the painting without meaningfully affecting other parts of the image, despite modifying all patches. This example is available at https://anonymous-saev.github.io/saev/demos/semseg?example=1633.
Figure 13: We find that feature DINOv2-24K/5876 fires for toilets, and that suppressing this feature removes the toilet without meaningfully affecting other parts of the image, despite modifying all patches. This example is available at https://anonymous-saev.github.io/saev/demos/semseg?example=1099.
Figure 14: After suppressing a “bed” feature (DINOv2-24K/18834), the segmentation head predicts “pillow” for the pillows and “table” for the bed spread. This example is available at https://anonymous-saev.github.io/saev/demos/semseg?example=1117.
Figure 15: We remove all cars from the scene by suppressing a “car”-like feature (DINOv2-24K/7235). Notably, the van on the right of the image is also removed. This example is available at https://anonymous-saev.github.io/saev/demos/semseg?example=1852.
12 Compute Requirements
We use a single NVIDIA A6000 48GB GPU for all our experiments.
Our experiments have several resource-intensive steps:
  1. 1.
    Record ViT activations on a large image dataset. This requires running ViT inference on many images and saving the activations to disk. For a ViT-B/16 on ImageNet-1K, this equates to 1.21.2M images $\times$ $196$ patches/image $\times$ $768$ floats/patch $\times$ 44 bytes/float = 722.5722.5GB. For a ViT-B/14 (DINOv2), the only difference is $256$ patches/image instead of $196$, for a total of 943.7943.7GB. This typically takes 6-12 hours depending on disk write speed.
  2. 2.
    Train SAEs on cached activations. This requires both good disk read speed and a GPU for acclerated training. Because SAEs are relatively small, our code enables parallel training of multiple SAEs at once to amortize the disk reads. We train 9 3232K-dimensional SAEs on a single A6000 GPU for 100100M patches in about 8 hours.
  3. 3.
    Save exemplar images for SAE features. This requires a GPU to run SAE inference on the training data and calculate the top-$k$ patches per feature and more disk space to save exemplar images. Saving 400 features with $k=128$ takes about 12 hours.